OpenStudy (anonymous):

(4g 2 - 9) ÷ (2g - 3)

OpenStudy (anonymous):

\[8g ^{3}+3g ^{2}-18g-9\]

OpenStudy (anonymous):

That not one of the choices here the choices 2g- 3 2g+ 3 2g+ 3 R 18

OpenStudy (anonymous):

(4g^2-9), that looks like a difference of squares, right? So how could you rewrite that expression differently?

OpenStudy (jackellyn):

When dividing variables with exponents, the exponents subtract 4g^2-9 ------- 2g-3 4/2=2 g^2/g=2-1=1=g^1=g -9/-3=3 2g+3